# Generalized Homogeneous Derivations on Graded Rings

### v5

@@ 1. Introduction
Derivations are fundamental mappings in ring theory that capture differential-like behavior while preserving the underlying algebraic structure. They play a central role in the study of ring-theoretic properties, especially in characterizing commutativity and detecting structural invariants in algebraic systems. Since their introduction, derivations have been the subject of numerous generalizations. A major development was Brešar’s introduction of generalized derivations [5], which opened a rich line of research (see, for example, [6, 7, 10, 12]. This broader framework has proved to be remarkably effective in extending many classical results that were originally established for standard derivations (cf. [4, 11, 9, 9].
@@
An ideal $\mathfrak{I}\subseteq R$ is graded when $\mathfrak{I}=\bigoplus_{\tau\in G}(\mathfrak{I}\cap R_{\tau})$, where we denote $\mathfrak{I}_{\tau}=\mathfrak{I}\cap R_{\tau}$. Ring homomorphisms $\varphi:R\rightarrow S$ between $G$-graded rings are graded if $\varphi(R_{\tau})\subseteq S_{\tau}$, $\operatorname{Hom}(R,S)^{gr}$ denoted the set of all graded homomorphisms. When $\mathfrak{I}$ is graded, quotient rings inherit the canonical grading structure $(R/\mathfrak{I})_{\tau}:=\{\bar{r}\in R/\mathfrak{I}\mid r\in R_{\tau}\}$. A graded ring $R$ is gr-prime if $aRb=\{0\}$ implies $a=0$ or $b=0$ for homogeneous elements $a,b\in\mathcal{H}(R)$, and gr-semiprime if $aRa=\{0\}$ implies $a=0$ for $a\in\mathcal{H}(R)$.
@@
An additive mapping $d:R\rightarrow R$ is a derivation if it satisfies the Leibniz rule $d(xy)=d(x)y+xd(y)$ for all $x,y\in R$. A derivation $d$ is homogeneous if $d(\mathcal{H}(R))\subseteq\mathcal{H}(R)$ ([8]). Inner derivations have the form $d(x)=[a,x]$ for some fixed $a\in R$, where [a,x]=axxa[a,x]=ax-xa denotes the commutator. A generalized derivation is an additive mapping $F:R\rightarrow R$ satisfying $F(xy)=F(x)y+xd(y)$ for all $x,y\in R$, where $d$ is the associated derivation of $F$.
@@
$R_{e}:=\left(\begin{array}[]{cccc}k&0&0&0\\ 0&k&0&0\\ 0&0&k&0\\ 0&0&0&k\end{array}\right),\hskip 17.07182ptR_{a}:=\left(\begin{array}[]{cccc}0&k&0&0\\ 0&0&k&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right),\hskip 17.07182ptR_{a^{2}}:=\left(\begin{array}[]{cccc}0&0&k&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{array}\right)$
$R_{a^{3}}:=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ k&0&0&0\\ 0&0&0&0\end{array}\right),\hskip 17.07182ptR_{b}:=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&k\\ 0&0&0&0\\ 0&k&0&0\end{array}\right),\hskip 17.07182ptR_{ab}:=\left(\begin{array}[]{cccc}0&0&0&k\\ 0&0&0&0\\ 0&0&0&0\\ k&0&0&0\end{array}\right)$
$R_{a^{4}b}:=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&k\\ 0&0&k&0\end{array}\right),\hskip 17.07182ptR_{a^{4}}:=\left(\begin{array}[]{cccc}0&0&0&0\\ k&0&0&0\\ 0&k&0&0\\ 0&0&0&0\end{array}\right),\hskip 17.07182ptR_{a^{2}b}=R_{a^{3}b}=\{0\}.$
@@
$R_{b},R_{a^{4}}]=R_{b}\circ R_{a^{4}}=\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&0&k\\ k&0&0&0\end{array}\right)\not\subset\mathcal{H}(R).$
@@ 2. Generalized homogeneous derivations
Let $R$ be a nontrivially $G$-graded ring. Then the following inclusions hold
@@
Corollary 2.1.
$|\mathfrak{Der}^{gh}_{G}(R)|$ is invariant under graded automorphisms.
@@
$\begin{array}[]{ccc}R&\stackrel{{\scriptstyle F_{R}}}{{\longrightarrow}}&R\\ \downarrow\phi&&\downarrow\phi\\ S&\stackrel{{\scriptstyle F_{S}}}{{\longrightarrow}}&S\end{array}\qquad\text{and}\qquad\begin{array}[]{ccc}R&\stackrel{{\scriptstyle d_{R}}}{{\longrightarrow}}&R\\ \downarrow\phi&&\downarrow\phi\\ S&\stackrel{{\scriptstyle d_{S}}}{{\longrightarrow}}&S\end{array}$
@@
Let $\{(R_{i},F_{i},d_{i})\}_{i\in I}$ be a finite family of objects in $\mathscr{G}_{G}^{h}$. By Proposition 2.4, we know that $R=\prod_{i\in I}R_{i}$ can be equipped with a generalized homogeneous derivation $(F,d)_{h}$ where $F((r_{i})_{i\in I})=(F_{i}(r_{i}))_{i\in I}$ and $d((r_{i})_{i\in I})=(d_{i}(r_{i}))_{i\in I}$, making $(R,F,d)$ an object in $\mathscr{G}_{G}^{h}$. To verify the universal property of products, let $(T,F_{T},d_{T})$ be an arbitrary object in $\mathscr{G}_{G}^{h}$ and let $\{\phi_{i}:T\rightarrow R_{i}\}_{i\in I}$ be a family of morphisms in $\mathscr{G}_{G}^{h}$. We must demonstrate the existence and uniqueness of a morphism $\phi:T\rightarrow R$ such that $\pi_{i}\circ\phi=\phi_{i}$ for each $i\in I$, where $\pi_{i}:R\rightarrow R_{i}$ denotes the canonical projection. Define $\phi:T\rightarrow R$ by $\phi(t)=(\phi_{i}(t))_{i\in I}$ for all $t\in T$. By construction, $\pi_{i}\circ\phi=\phi_{i}$ for each $i\in I$. To verify that $\phi$ is a morphism in $\mathscr{G}_{G}^{h}$, we confirm its compatibility with the generalized homogeneous derivations: For any $t\in T$
$\phi(F_{T}(t))$ $=(\phi_{i}(F_{T}(t)))_{i\in I}$
$=(F_{i}(\phi_{i}(t)))_{i\in I}$
$=F((\phi_{i}(t))_{i\in I})$
$=F(\phi(t)).$
Similarly, for the derivation component
$\phi(d_{T}(t))$ $=(\phi_{i}(d_{T}(t)))_{i\in I}$
$=(d_{i}(\phi_{i}(t)))_{i\in I}$
$=d((\phi_{i}(t))_{i\in I})$
$=d(\phi(t)).$
The uniqueness of $\phi$ follows directly from the universal property of the categorical product. Indeed, if $\psi:T\rightarrow R$ is another morphism in $\mathscr{G}_{G}^{h}$ such that $\pi_{i}\circ\psi=\phi_{i}$ for each $i\in I$, then for any $t\in T$
$\psi(t)=(\pi_{i}(\psi(t)))_{i\in I}=(\phi_{i}(t))_{i\in I}=\phi(t).$
Hence, $\psi=\phi$. ∎
@@
Let $H\in\mathfrak{Der}^{h}_{G}(R)\cap\mathcal{C}_{G}(R)$. The derivation property gives $H(xy)=H(x)y+xH(y)$, while the condition $H\in\mathcal{C}_{G}(R)$ implies $H(xy)=H(x)y$. Hence $xH(y)=0$ for all $x,y\in R$. Thus $xRH(y)=\{0\}$ for all $x,y\in R$. For any $r\in\mathcal{H}(R)\setminus\{0\}$, we obtain $rRH(y)=\{0\}$. By [1, Proposition 2.1], it follows that $H(y)=0$ for all $y\in R$. Therefore $H\equiv 0$. ∎
@@
That $(F^{\prime},d^{\prime})_{h}\in p\mathfrak{Der}^{gh}_{G}(R[t])$ follows from direct computations. For homogeneity, observe that if $f(t)=\sum_{i=0}^{n}r_{i}t^{i}\in R[t]_{\tau}$, then $r_{i}\in R_{\tau}$, so $F(r_{i}),d(r_{i})\in R_{\tau}$ by the homogeneity of $(F,d)_{h}$. Thus $F^{\prime}(f),d^{\prime}(f)\in R[t]_{\tau}$. For injectivity, if $(F,d)_{h}\neq(0,0)_{h}$, then there exists $r\in R$ such that either $F(r)\neq 0$ or $d(r)\neq 0$, which implies $F^{\prime}(r)\neq 0$ or $d^{\prime}(r)\neq 0$ when $r$ is viewed as a constant polynomial. ∎
@@
We extend the definition of $F_{R\otimes S}$ and $d_{R\otimes S}$ to all of $R\otimes_{k}S$ by linearity. To verify that $(F_{R\otimes S},d_{R\otimes S})_{h}$ is a generalized homogeneous derivation, we must verify that
$F_{R\otimes S}(uv)=F_{R\otimes S}(u)v+ud_{R\otimes S}(v),\hskip 8.5359ptd_{R\otimes S}(uv)=d_{R\otimes S}(r_{1}r_{2}\otimes s_{1}s_{2})$
for all $u,v\in R\otimes_{K}S$. By linearity, it suffices to check this identity for homogeneous tensors $u=r_{1}\otimes s_{1}$ and $v=r_{2}\otimes s_{2}$. We have
$F_{R\otimes S}((r_{1}\otimes s_{1})(r_{2}\otimes s_{2}))$ $=F_{R\otimes S}(r_{1}r_{2}\otimes s_{1}s_{2})$
$=F_{R}(r_{1}r_{2})\otimes s_{1}s_{2}+r_{1}r_{2}\otimes F_{S}(s_{1}s_{2})$
@@
$=F_{R}(r_{1})r_{2}\otimes s_{1}s_{2}+r_{1}d_{R}(r_{2})\otimes s_{1}s_{2}+r_{1}r_{2}\otimes F_{S}(s_{1})s_{2}+r_{1}r_{2}\otimes s_{1}d_{S}(s_{2})$
$=[F_{R}(r_{1})\otimes s_{1})+r_{1}\otimes F_{S}(s_{1})](r_{2}\otimes s_{2})+(r_{1}\otimes s_{1})[d_{R}(r_{2})\otimes s_{2}+r_{2}\otimes d_{S}(s_{2})]$
$=F_{R\otimes S}(r_{1}\otimes s_{1})(r_{2}\otimes s_{2})+(r_{1}\otimes s_{1})d_{R\otimes S}(r_{2}\otimes s_{2})$
For the homogeneity condition, observe that if $r\in R_{\tau}$ and $s\in S_{\sigma}$, then $r\otimes s\in(R\otimes_{K}S)_{\tau\sigma}$. Since $F_{R}(r)\in R_{\tau}$ and $F_{S}(s)\in S_{\sigma}$, we have $F_{R\otimes S}(r\otimes s)=F_{R}(r)\otimes s+r\otimes F_{S}(s)\in(R\otimes_{K}S)_{\tau\sigma}$ Additionally, we must verify that $d_{R\otimes S}$ satisfies the Leibniz rule for a homogeneous derivation. For homogeneous tensors $u=r_{1}\otimes s_{1}$ and $v=r_{2}\otimes s_{2}$
$d_{R\otimes S}(uv)$ $=d_{R\otimes S}(r_{1}r_{2}\otimes s_{1}s_{2})$
$=(d_{R}(r_{1})r_{2}+r_{1}d_{R}(r_{2}))\otimes s_{1}s_{2}+r_{1}r_{2}\otimes(d_{S}(s_{1})s_{2}+s_{1}d_{S}(s_{2}))$
$=d_{R}(r_{1})r_{2}\otimes s_{1}s_{2}+r_{1}d_{R}(r_{2})\otimes s_{1}s_{2}+r_{1}r_{2}\otimes d_{S}(s_{1})s_{2}+r_{1}r_{2}\otimes s_{1}d_{S}(s_{2})$
$=d_{R}(r_{1})\otimes s_{1}(r_{2}\otimes s_{2})+r_{1}\otimes d_{S}(s_{1})(r_{2}\otimes s_{2})$
$=d_{R\otimes S}(r_{1}\otimes s_{1})(r_{2}\otimes s_{2})+(r_{1}\otimes s_{1})d_{R\otimes S}(r_{2}\otimes s_{2})$
Moreover, $d_{R\otimes S}(r\otimes s)\in(R\otimes_{K}S)_{\tau\sigma}$. Hence, $(F_{R\otimes S},d_{R\otimes S})_{h}\in p\mathfrak{Der}^{gh}_{G}(R\otimes_{K}S)$. ∎
@@ 4.1. Results on homogeneous derivations
  • (1)
    Let $a=\sum_{g\in G}a_{\tau}\in R$ and $b\in\mathcal{H}(R)\setminus\{0\}$ such that $a\mathfrak{I}b=\{0\}$. Then, for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ we have $arb=0$, which implies $\sum_{\tau\in G}a_{\tau}rb=0$. Since every element of $R$ has a unique homogeneous decomposition, it follows that $a_{\tau}rb=0$ for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and all $\tau\in G$. Hence $a_{\tau}\mathfrak{I}Rb=\{0\}$ for all $\tau\in G$. By [1, Proposition 2.1], we obtain $a_{\tau}\mathfrak{I}=\{0\}$ for all $\tau\in G$, i.e. $a_{\tau}R\mathfrak{I}=\{0\}$ for all $\tau\in G$. Thus $a_{\tau}=0$ for all $\tau\in G$, and therefore $a=0$.
  • (2)
    Suppose that $ad(x)=0$ and $a\neq 0$. Replacing $x$ by $xy$, we obtain $axd(y)=0$ for all $x,y\in R$. In particular, $aRd(x)=\{0\}$ for all $x\in R$. Hence $aRd(r)=\{0\}$ for all $r\in\mathcal{H}(R)$. By [1, Proposition 2.1], we conclude that $d(r)=0$ for all $r\in\mathcal{H}(R)$, and thus $d=0$.
@@
(1) $[z,[x,y]]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. Replacing $y$ by $yx$ in (1) and simplifying, we obtain
(2) $[x,y][z,x]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. Substituting $zy$ for $z$ in (2) yields
$[x,y]z[x,y]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$, which implies
$x,y]R[x,y]=\{0\}$
for all $x,y\in\mathfrak{I}$. Since $\mathfrak{I}$ is a graded ideal, we also have
$r_{1},r_{2}]R[r_{1},r_{2}]=\{0\}$
for all $r_{1},r_{2}\in\mathfrak{I}\cap\mathcal{H}(R)$. By gr-primeness of $R$, it follows that $r_{1},r_{2}]=0$ for all $r_{1},r_{2}\in\mathfrak{I}\cap\mathcal{H}(R)$. Hence $[x,y]=0$ for all $x,y\in\mathfrak{I}$, so $\mathfrak{I}$ is commutative. In view of [1, Proposition 2.1], $R$ is then commutative.
Now assume that $x\circ y\in Z(R)$ for all $x,y\in\mathfrak{I}$. Then
(3) $x\circ y,z]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. Replacing $y$ by $yx$ in (3) and simplifying, we obtain
(4) $(x\circ y)[x,z]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. Substituting $sz$ for $z$ in (4), we get
$(x\circ y)s[x,z]=0$
for all $x,y\in\mathfrak{I}$ and $s,z\in R$. Hence
$(x\circ y)R[x,z]=\{0\}$
for all $x,y\in\mathfrak{I}$ and $z\in R$. In particular,
$(r_{1}\circ r_{2})R[r_{1},z]=\{0\}$
for all $r_{1},r_{2}\in\mathfrak{I}\cap\mathcal{H}(R)$ and $z\in R$. According to [1, Proposition 2.1], it follows that
$r_{1}\circ r_{2}=0\quad\text{or}\quad[r_{1},z]=0$
for all $r_{1},r_{2}\in\mathfrak{I}\cap\mathcal{H}(R)$ and $z\in R$. Thus
$x\circ y=0\quad\text{or}\quad[x,z]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. In the latter case, $\mathfrak{I}$ is a central graded ideal, and applying [1, Proposition 2.1], we conclude that $R$ is commutative.
We may therefore assume that $x\circ y=0$ for all $x,y\in\mathfrak{I}$. Replacing $y$ by $yz$ gives
$y[x,z]=0$
for all $x,y\in\mathfrak{I}$ and $z\in R$. Since $\mathfrak{I}$ is a nonzero ideal of $R$, there exists $a\in\mathfrak{I}\setminus\{0\}$ such that
$a[x,z]=0$
for all $x\in\mathfrak{I}$ and $z\in R$. Fix $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and let $d_{r}$ be the inner homogeneous derivation associated with $r$, i.e. $d_{r}(z)=[r,z]$ for $z\in R$. Then
$ad_{r}(z)=0$
for all $z\in R$. By Lemma 4.1(2), we obtain $d_{r}(z)=[r,z]=0$ for all $z\in R$. Hence $[x,z]=0$ for all $x\in\mathfrak{I}$ and $z\in R$. In both cases, we find that $\mathfrak{I}$ is a central graded ideal of $R$. Therefore, $R$ is commutative. ∎
@@
(5) $d_{1}d_{2}(x)\in Z(R)$
for all $x\in R$. Replacing $x$ by $[x,y]$ in (5) and expanding, we obtain
(6) $d_{2}(x),d_{1}(y)]+[d_{1}(x),d_{2}(y)]\in Z(R)$
for all $x,y\in R$. Putting $y=d_{2}(z)$ in (6) yields
$d_{1}(x),d_{2}^{2}(z)]\in Z(R)$
for all $x,z\in R$. In particular,
$d_{2}^{2}(r),d_{1}(y)]\in Z(R)$
for all $r\in\mathcal{H}(R)$ and $y\in R$. By [1, Lemma 2.2], it follows that either $d_{2}^{2}(r)\in Z(R)$ for all $r\in\mathcal{H}(R)$ or $d_{1}=0$. The latter is impossible by assumption, so $d_{2}^{2}(x)\in Z(R)$ for all $x\in R$. Taking $[x,z]$ instead of $x$, we obtain
$2[d_{2}(x),d_{2}(z)]\in Z(R)$
for all $x,z\in R$. Since $\operatorname{char}R\neq 2$, it follows that
$d_{2}(x),d_{2}(z)]\in Z(R)$
for all $x,z\in R$. By [1, Theorem 3.5], we conclude that $R$ is commutative. ∎
The following example proves that the gr-primeness hypothesis in Theorem 4.1 is not superfluous. In particular, our theorem cannot be extended to gr-semiprime rings.
@@
$\begin{array}[]{ccc}d_{1}\left(f,M\right)=\left(t_{2}t_{4}\frac{\partial f}{\partial t_{1}},0\right),&&d_{2}\left(f,M\right)=\left(t_{1}t_{3}\frac{\partial f}{\partial t_{2}},0\right)\end{array}.$

### v6

@@ Abstract.
Key words and phrases:
Generalized homogeneous derivations, graded rings, gr-prime rings, gr-semiprime rings, commutativity, graded modules
2020 Mathematics Subject Classification:
16W25, 16W50, 13A02, 16N60
@@ 1. Introduction
Derivations are fundamental mappings in ring theory that capture differential-like behavior while preserving the underlying algebraic structure. They play a central role in the study of ring-theoretic properties, especially in characterizing commutativity and detecting structural invariants in algebraic systems. Since their introduction, derivations have been the subject of numerous generalizations. A major development was Brešar’s introduction of generalized derivations [5], which opened a rich line of research (see, for example, [6, 7, 10, 12]). This broader framework has proved to be remarkably effective in extending many classical results that were originally established for standard derivations (cf.[4, 11, 9, 9]).
@@
An ideal $\mathfrak{I}\subseteq R$ is graded when $\mathfrak{I}=\bigoplus_{\tau\in G}(\mathfrak{I}\cap R_{\tau})$, where we denote $\mathfrak{I}_{\tau}=\mathfrak{I}\cap R_{\tau}$. Ring homomorphisms $\varphi:R\rightarrow S$ between $G$-graded rings are graded if $\varphi(R_{\tau})\subseteq S_{\tau}$; $\operatorname{Hom}(R,S)^{gr}$ denotes the set of all graded homomorphisms. When $\mathfrak{I}$ is graded, quotient rings inherit the canonical grading structure $(R/\mathfrak{I})_{\tau}:=\{\bar{r}\in R/\mathfrak{I}\mid r\in R_{\tau}\}$. A graded ring $R$ is gr-prime if $aRb=\{0\}$ implies $a=0$ or $b=0$ for homogeneous elements $a,b\in\mathcal{H}(R)$, and gr-semiprime if $aRa=\{0\}$ implies $a=0$ for $a\in\mathcal{H}(R)$.
@@
An additive mapping $d:R\rightarrow R$ is a derivation if it satisfies the Leibniz rule $d(xy)=d(x)y+xd(y)$ for all $x,y\in R$. A derivation $d$ is homogeneous if $d(\mathcal{H}(R))\subseteq\mathcal{H}(R)$ [8]. Inner derivations have the form $d(x)=[a,x]$ for some fixed $a\in R$, where [a,x]=axxa[a,x]=ax-xa denotes the commutator. A generalized derivation is an additive mapping $F:R\rightarrow R$ satisfying $F(xy)=F(x)y+xd(y)$ for all $x,y\in R$, where $d$ is the associated derivation of $F$.
@@
$R_{e}:=\begin{pmatrix}k&0&0&0\\ 0&k&0&0\\ 0&0&k&0\\ 0&0&0&k\end{pmatrix},\quad R_{a}:=\begin{pmatrix}0&k&0&0\\ 0&0&k&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix},\quad R_{a^{2}}:=\begin{pmatrix}0&0&k&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}$,
$R_{a^{3}}:=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ k&0&0&0\\ 0&0&0&0\end{pmatrix},\quad R_{b}:=\begin{pmatrix}0&0&0&0\\ 0&0&0&k\\ 0&0&0&0\\ 0&k&0&0\end{pmatrix},\quad R_{ab}:=\begin{pmatrix}0&0&0&k\\ 0&0&0&0\\ 0&0&0&0\\ k&0&0&0\end{pmatrix}$,
$R_{a^{4}b}:=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&k\\ 0&0&k&0\end{pmatrix},\quad R_{a^{4}}:=\begin{pmatrix}0&0&0&0\\ k&0&0&0\\ 0&k&0&0\\ 0&0&0&0\end{pmatrix},\quad R_{a^{2}b}=R_{a^{3}b}=\{0\}.$
@@
$R_{b},R_{a^{4}}]=R_{b}\circ R_{a^{4}}=\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&0&k\\ k&0&0&0\end{pmatrix}\not\subset\mathcal{H}(R).$
@@ 2. Generalized homogeneous derivations
Let $R$ be a nontrivially $G$-graded ring. Then the following inclusions hold:
@@
[…]
@@
Let $\{(R_{i},F_{i},d_{i})\}_{i\in I}$ be a finite family of objects in $\mathscr{G}_{G}^{h}$. By Proposition 2.10, $R=\prod_{i\in I}R_{i}$ can be equipped with a generalized homogeneous derivation $(F,d)_{h}$ where $F((r_{i})_{i\in I})=(F_{i}(r_{i}))_{i\in I}$ and $d((r_{i})_{i\in I})=(d_{i}(r_{i}))_{i\in I}$, making $(R,F,d)$ an object in $\mathscr{G}_{G}^{h}$. To verify the universal property of products, let $(T,F_{T},d_{T})$ be an arbitrary object in $\mathscr{G}_{G}^{h}$ and let $\{\phi_{i}:T\rightarrow R_{i}\}_{i\in I}$ be a family of morphisms in $\mathscr{G}_{G}^{h}$. Define $\phi:T\rightarrow R$ by $\phi(t)=(\phi_{i}(t))_{i\in I}$ for all $t\in T$. By construction, $\pi_{i}\circ\phi=\phi_{i}$ for each $i\in I$. For any $t\in T$:
$\phi(F_{T}(t))$ $=(\phi_{i}(F_{T}(t)))_{i\in I}=(F_{i}(\phi_{i}(t)))_{i\in I}=F((\phi_{i}(t))_{i\in I})=F(\phi(t)),$
and similarly $\phi(d_{T}(t))=d(\phi(t))$. Uniqueness of $\phi$ follows from the universal property of the categorical product. ∎
@@
Let $H\in\mathfrak{Der}^{h}_{G}(R)\cap\mathcal{C}_{G}(R)$. The derivation property gives $H(xy)=H(x)y+xH(y)$, while the condition $H\in\mathcal{C}_{G}(R)$ implies $H(xy)=H(x)y$. Hence $xH(y)=0$ for all $x,y\in R$, so $xRH(y)=\{0\}$ for all $x,y\in R$. For any $r\in\mathcal{H}(R)\setminus\{0\}$, we obtain $rRH(y)=\{0\}$. By [1, Proposition 2.1], it follows that $H(y)=0$ for all $y\in R$, hence $H\equiv 0$. ∎
@@
That $(F^{\prime},d^{\prime})_{h}\in p\mathfrak{Der}^{gh}_{G}(R[t])$ follows from direct computations. For homogeneity, observe that if $f(t)=\sum_{i=0}^{n}r_{i}t^{i}\in R[t]_{\tau}$, then $r_{i}\in R_{\tau}$, so $F(r_{i}),d(r_{i})\in R_{\tau}$. Thus $F^{\prime}(f),d^{\prime}(f)\in R[t]_{\tau}$. For injectivity, if $(F,d)_{h}\neq(0,0)_{h}$, then there exists $r\in R$ such that $F(r)\neq 0$ or $d(r)\neq 0$, which implies $F^{\prime}(r)\neq 0$ or $d^{\prime}(r)\neq 0$ when $r$ is viewed as a constant polynomial. ∎
@@
We extend the definition of $F_{R\otimes S}$ and $d_{R\otimes S}$ to all of $R\otimes_{k}S$ by linearity. By linearity, it suffices to check the identity $F_{R\otimes S}(uv)=F_{R\otimes S}(u)v+ud_{R\otimes S}(v)$ for homogeneous tensors $u=r_{1}\otimes s_{1}$ and $v=r_{2}\otimes s_{2}$:
$F_{R\otimes S}((r_{1}\otimes s_{1})(r_{2}\otimes s_{2}))$ $=F_{R}(r_{1}r_{2})\otimes s_{1}s_{2}+r_{1}r_{2}\otimes F_{S}(s_{1}s_{2})$
@@
$=(F_{R}(r_{1})\otimes s_{1})(r_{2}\otimes s_{2})+(r_{1}\otimes s_{1})(d_{R}(r_{2})\otimes s_{2}+r_{2}\otimes d_{S}(s_{2}))$
$=F_{R\otimes S}(r_{1}\otimes s_{1})(r_{2}\otimes s_{2})+(r_{1}\otimes s_{1})d_{R\otimes S}(r_{2}\otimes s_{2}).$
For homogeneity, if $r\in R_{\tau}$ and $s\in S_{\sigma}$, then $r\otimes s\in(R\otimes_{k}S)_{\tau\sigma}$. Since $F_{R}(r)\in R_{\tau}$ and $F_{S}(s)\in S_{\sigma}$, we have $F_{R\otimes S}(r\otimes s)\in(R\otimes_{k}S)_{\tau\sigma}$. The verification that $d_{R\otimes S}$ satisfies the Leibniz rule is analogous. Hence $(F_{R\otimes S},d_{R\otimes S})_{h}\in p\mathfrak{Der}^{gh}_{G}(R\otimes_{k}S)$. ∎
@@ 4.1. Results on homogeneous derivations
(1) Let $a=\sum_{g\in G}a_{\tau}\in R$ and $b\in\mathcal{H}(R)\setminus\{0\}$ such that $a\mathfrak{I}b=\{0\}$. Then, for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ we have $arb=0$, which implies $\sum_{\tau\in G}a_{\tau}rb=0$. Since every element of $R$ has a unique homogeneous decomposition, it follows that $a_{\tau}rb=0$ for all $r\in\mathfrak{I}\cap\mathcal{H}(R)$ and all $\tau\in G$. Hence $a_{\tau}\mathfrak{I}Rb=\{0\}$ for all $\tau\in G$. By [1, Proposition 2.1], we obtain $a_{\tau}\mathfrak{I}=\{0\}$ for all $\tau\in G$, i.e., $a_{\tau}R\mathfrak{I}=\{0\}$ for all $\tau\in G$. Thus $a_{\tau}=0$ for all $\tau\in G$, and therefore $a=0$.
(2) Suppose that $ad(x)=0$ and $a\neq 0$. Replacing $x$ by $xy$, we obtain $axd(y)=0$ for all $x,y\in R$. In particular, $aRd(x)=\{0\}$ for all $x\in R$. Hence $aRd(r)=\{0\}$ for all $r\in\mathcal{H}(R)$. By [1, Proposition 2.1], we conclude that $d(r)=0$ for all $r\in\mathcal{H}(R)$, and thus $d=0$. ∎
@@
$[z,[x,y]]=0$ (4.1)
for all $x,y\in\mathfrak{I}$ and $z\in R$. Replacing $y$ by $yx$ in (4.1) and simplifying, we obtain
$[x,y][z,x]=0$ (4.2)
for all $x,y\in\mathfrak{I}$ and $z\in R$. Substituting $zy$ for $z$ in (4.2) yields $[x,y]z[x,y]=0$ for all $x,y\in\mathfrak{I}$ and $z\in R$, which implies $x,y]R[x,y]=\{0\}$ for all $x,y\in\mathfrak{I}$. Since $\mathfrak{I}$ is a graded ideal, $r_{1},r_{2}]R[r_{1},r_{2}]=\{0\}$ for all $r_{1},r_{2}\in\mathfrak{I}\cap\mathcal{H}(R)$. By gr-primeness of $R$, it follows that $r_{1},r_{2}]=0$, hence $[x,y]=0$ for all $x,y\in\mathfrak{I}$, so $\mathfrak{I}$ is commutative. In view of [1, Proposition 2.1], $R$ is then commutative.
Now assume that $x\circ y\in Z(R)$ for all $x,y\in\mathfrak{I}$. Then $x\circ y,z]=0$ for all $x,y\in\mathfrak{I}$ and $z\in R$. Replacing $y$ by $yx$ and simplifying:
$(x\circ y)[x,z]=0$ (4.3)
for all $x,y\in\mathfrak{I}$ and $z\in R$. Substituting $sz$ for $z$, we get $(r_{1}\circ r_{2})R[r_{1},z]=\{0\}$ for all $r_{1},r_{2}\in\mathfrak{I}\cap\mathcal{H}(R)$ and $z\in R$. By [1, Proposition 2.1], either $x\circ y=0$ or $[x,z]=0$ for all $x,y\in\mathfrak{I}$ and $z\in R$. In the latter case $\mathfrak{I}$ is a central graded ideal and $R$ is commutative. We may therefore assume $x\circ y=0$ for all $x,y\in\mathfrak{I}$. Replacing $y$ by $yz$ gives $y[x,z]=0$, and arguing as in Lemma 4.2(2), we find that $[x,z]=0$ for all $x\in\mathfrak{I}$ and $z\in R$. In both cases $\mathfrak{I}$ is a central graded ideal, so $R$ is commutative. ∎
@@
$d_{1}d_{2}(x)\in Z(R)$ (4.4)
for all $x\in R$. Replacing $x$ by $[x,y]$ in (4.4) and expanding, we obtain
$d_{2}(x),d_{1}(y)]+[d_{1}(x),d_{2}(y)]\in Z(R)$ (4.5)
for all $x,y\in R$. Putting $y=d_{2}(z)$ in (4.5) yields $d_{1}(x),d_{2}^{2}(z)]\in Z(R)$ for all $x,z\in R$. In particular, $d_{2}^{2}(r),d_{1}(y)]\in Z(R)$ for all $r\in\mathcal{H}(R)$ and $y\in R$. By [1, Lemma 2.2], either $d_{2}^{2}(r)\in Z(R)$ for all $r\in\mathcal{H}(R)$ or $d_{1}=0$. The latter is impossible by assumption, so $d_{2}^{2}(x)\in Z(R)$ for all $x\in R$. Taking $[x,z]$ instead of $x$, we obtain $2[d_{2}(x),d_{2}(z)]\in Z(R)$ for all $x,z\in R$. Since $\operatorname{char}R\neq 2$, it follows that $d_{2}(x),d_{2}(z)]\in Z(R)$ for all $x,z\in R$. By [1, Theorem 3.5], we conclude that $R$ is commutative. ∎
The following example proves that the gr-primeness hypothesis in Theorem 4.3 is not superfluous; in particular, the theorem cannot be extended to gr-semiprime rings.
@@
$d_{1}\left(f,M\right)=\left(t_{2}t_{4}\frac{\partial f}{\partial t_{1}},0\right),\qquad d_{2}\left(f,M\right)=\left(t_{1}t_{3}\frac{\partial f}{\partial t_{2}},0\right).$